# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v2

@@ A.4 More Comparison on HumanEval for Code Generation
MoRA 47.6 60.9 70.1
A.5 performance robustness
MoRA employs a sparse mixture of previously learned and newly introduced rank-1 experts to capture both shared and task-specific knowledge, resulting in substantially improved Last performance. To assess statistical significance and robustness, we report mean and standard deviation over three independent runs (Table 13). MoRA consistently outperforms competing methods across all metrics and exhibits lower variance, highlighting its effectiveness and stability in continual-learning scenarios.
Table 13: Comparison to InfLoRA and performance robustness. We report mean and standard deviation across 3 independent runs. Best performances are marked in bold.
Method
Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
@@
   MoRA $\mathbf{74.5^{\pm 0.51}}$ $\mathbf{38.1^{\pm 0.24}}$ $\mathbf{46.9^{\pm 0.56}}$ $65.3^{\pm 0.44}$ $82.9^{\pm 0.18}$ $\mathbf{45.8^{\pm 0.31}}$ $88.2^{\pm 0.15}$ $\mathbf{65.1^{\pm 0.35}}$ $62.9^{\pm 0.10}$ $\mathbf{63.3^{\pm 0.26}}$
@@
   MoRA $\mathbf{44.1^{\pm 0.24}}$ $\mathbf{81.6^{\pm 0.34}}$ $\mathbf{64.6^{\pm 0.34}}$ $\mathbf{79.6^{\pm 0.37}}$ $\mathbf{83.9^{\pm 0.36}}$ $84.4^{\pm 0.15}$ $\mathbf{66.5^{\pm 0.24}}$ $89.7^{\pm 0.07}$ $\mathbf{68.4^{\pm 0.38}}$ $64.1^{\pm 0.09}$ $\mathbf{72.7^{\pm 0.17}}$
@@
   MoRA $\mathbf{37.7^{\pm 0.28}}$ $81.5^{\pm 0.22}$ $\mathbf{70.7^{\pm 0.49}}$ $\mathbf{92.4^{\pm 0.20}}$ $\mathbf{95^{\pm 0.34}}$ $\mathbf{86^{\pm 0.13}}$ $\mathbf{97.6^{\pm 0.19}}$ $92.6^{\pm 0.10}$ $\mathbf{81^{\pm 0.35}}$ $\mathbf{74.7^{\pm 0.06}}$ $\mathbf{80.9^{\pm 0.12}}$
A.6 More comparisons on X-TAIL
To further validate MoRA’s robustness, we evaluated it under other continual-learning task orderings, i.e., X-TAIL (Order 2), as shown in Table 14. The results align with those in Table 1, confirming that MoRA consistently achieves state-of-the-art performance.
@@ A.6 More comparisons on X-TAIL
Method
Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
@@
   Zero-shot (Radford et al., 2021) 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
   LwF (Li & Hoiem, 2017) 20.0 74.1 79.6 58.1 34.1 48.9 27.7 64.4 15.1 46.9
   WiSE-FT (Wortsman et al., 2022) 21.3 79.5 83.3 61.0 39.9 56.5 29.6 68.0 20.8 51.1
   ZSCL (Zheng et al., 2023) 23.0 84.3 87.2 63.0 42.1 65.2 34.6 71.4 40.9 56.9
   MoE-Adapter (Yu et al., 2024) 17.1 87.2 87.5 58.4 12.6 65.5 35.9 70.0 17.9 50.2
   RAIL-Primal (Xu et al., 2024) 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
   CoDyRA (Lu et al., 2024) 23.6 89.2 83 62 51 71.4 38 77.4 39 59.4
   MoRA 23.6 88.7 83.4 62.6 51.2 69.9 39.3 77.5 39 59.5
@@
   LwF (Li & Hoiem, 2017) 49.0 27.4 69.7 83.0 65.7 42.2 63.5 33.1 68.5 17.5 52.0
   WiSE-FT (Wortsman et al., 2022) 57.9 29.6 77.8 85.4 68.0 51.6 69.3 35.5 71.0 23.0 56.9
   ZSCL (Zheng et al., 2023) 74.4 36.4 86.7 88.7 68.9 50.0 75.1 40.1 72.5 43.7 63.6
   MoE-Adapter (Yu et al., 2024) 74.4 38.6 87.7 87.3 67.9 50.6 76.5 43.7 72.3 18.8 61.8
   RAIL-Primal (Xu et al., 2024) 77.9 40.4 85.6 83.3 68.3 62.2 76.6 45.8 80.4 41.7 66.2
   CoDyRA (Lu et al., 2024) 80 39.2 92.5 85.2 69.2 73.7 79.6 46.2 78.6 44.1 68.8
   MoRA 80.2 40.1 92.5 84.7 70.1 74 80.1 48.7 78.4 44.4 69.3
@@
   LwF (Li & Hoiem, 2017) 29.6 17.5 63.0 83.8 67.7 44.9 79.3 44.8 84.6 39.0 55.4
   WiSE-FT (Wortsman et al., 2022) 46.1 23.5 71.3 85.7 70.2 59.1 85.5 47.9 82.4 42.8 61.5
   ZSCL (Zheng et al., 2023) 71.7 35.3 86.5 89.2 71.8 52.3 89.8 52.0 77.1 68.4 69.4
   MoE-Adapter (Yu et al., 2024) 75.1 41.1 87.9 87.1 74.1 89.7 92.6 61.2 81.0 27.4 71.7
   RAIL-Primal (Xu et al., 2024) 77.7 41.9 86.1 83.3 71.8 91.6 97.3 66.4 94.8 86.9 79.8
   CoDyRA (Lu et al., 2024) 79 38.6 92.6 86.4 74.7 95.2 93 64.7 81.9 92.2 79.8
   MoRA 79.3 38.9 93.1 85.4 74.9 96.4 94.1 69.9 82 92.9 80.7
@@
LWF (Li & Hoiem, 2017) 129.6 32172
ZSCL (Zheng et al., 2023) 129.6 26290
MoE-Adapters (Yu et al., 2024) 59.8 22358
CoDyRA (Lu et al., 2024) 4.4 21770
MoRA 4.4 21090
A.7 Computation Cost
Table 10 summarizes the per‐task trainable parameters of various continual‐learning methods. Standard LoRA (Hu et al., 2021), CoDyRA (Lu et al., 2024), and O-LoRA (Wang et al., 2023) each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix. Mixture‐of‐Experts variants such as MoE-LoRA and MoE-Adapter (Yu et al., 2024) additionally train a router module to controll the usage of each LoRA experts, inducing $d_{\text{in}}$ additional paramters for each experts. LB-CL (Qiao & Mahdavi, 2024) introduce $r$ additional parameters, mimic the singular values of SVD.
By contrast, MoRA requires only $rd_{\text{in}}+kd_{\text{out}}$ activated trainable parameters per task, where $k\leq r$ is the activation budget, and the trainable parameter is at most the same as a standard LoRA. Despite the small number of parameter activated and trained, MoRA achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one‐third of the activated parameters of a standard LoRA.
@@ A.7 Computation Cost
LWF (Li & Hoiem, 2017) and ZSCL (Zheng et al., 2023) perform full‐parameter fine‐tuning, consuming the most parameters and memory. MoE‐Adapters (Yu et al., 2024) maintains a router with 22 rank‐64 adapter experts (top-2 activated) and a DDAS domain predictor. CoDyRA (Lu et al., 2024) trains a single rank‐16 LoRA per task, reducing its footprint to 4.4 M parameters. MoRA introduces 16 rank-1 experts per task, with no additional router, for a total of 4.4 M trainable parameters and keeps a low gpu memory usage, thanks to our novel self‐activated sparse mixture of ranks design.
@@
(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
@@
(b) Rank Activations of MoRA on data from Task 1 after learning Task 2.
@@
(c) Rank Activations of MoRA on data from Task 1 after learning Task 10.
@@
(a) Rank Activations of MoRA on data from Task 1 after learning Task 1.
@@
(b) Rank Activations of MoRA on data from Task 9 after learning Task 1.
@@
(c) Rank Activations of MoRA on data from Task 9 after learning Task 9.
@@
A.8 Extended Visualizations of Rank Activations
@@
MoRA retains task‐specific semantics without forgetting. Fig. 4 shows the activation maps for the same Task 1 image after training on Task 1 (a), Task 2 (b), and the last Task 10 (c). Patches corresponding to the airplane object are outlined in orange. In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRA has effectively memorized the airplane‐specific knowledge into Rank 0. Even after 10 subsequent tasks, this pattern remains unchanged, indicating that later updates do not overwrite or interfere with the learned airplane representations. In other words, MoRA effectively memorizes and preserves task‐relevant semantics, thereby mitigating catastrophic forgetting.
MoRA encodes generic semantics that are reused across tasks. Fig. 5 examines an input image from Task 9 before and after learning Task 9. Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky‐background patches, demonstrating that MoRA has stored a generic “blue sky” concept in this rank. In panel (b), when we infer on the Task 9 image before training on Task 9, Rank 11 is again activated for the sky regions, confirming that MoRA reuses this shared knowledge for unseen data. Finally, panel (c) shows the activation map after learning Task 9: Rank 11 remains dedicated to the sky background, while newly initialized ranks specialize in the “car” object semantics. This persistent reuse of Rank 11 across tasks illustrates MoRA’s ability to capture and retain common features as reusable memory slots, reducing redundancy and facilitating knowledge reuse.
Figure 6: Statistical analyses on the number of ranks required to capture 99% of cumulative sum (indicated in red dashed line) of all rank activations. Activations were gathered from the model after training on all tasks, and results are shown for a representative selection of layers and positions within the pre-trained model.
Figure 7: Statistical analyses on the number of ranks required to capture 99% of cumulative sum (indicated in red dashed line) of all rank activations. Activations were gathered from the model after training on all tasks, and results are shown for a representative selection of layers and positions within the pre-trained model.
@@
A.9 Statistical Analyses of Contributing Rank Activations

### v3

@@ A.7 More Comparison on HumanEval for Code Generation
MoRAM 47.6 60.9 70.1
A.8 Robustness
MoRAM employs a sparse mixture of previously learned and newly introduced rank-1 experts to capture both shared and task-specific knowledge, resulting in substantially improved Last performance. To assess statistical significance and robustness, we report mean and standard deviation over three independent runs (Table 16). MoRAM consistently outperforms competing methods across all metrics and exhibits lower variance, highlighting its effectiveness and stability in continual-learning scenarios.
Table 16: Comparison to InfLoRA and performance robustness. We report the mean and standard deviation across 3 independent runs. Best performances are marked in bold.
      Method Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
@@ A.8 Robustness
     MoRAM $\mathbf{74.5^{\pm 0.51}}$ $\mathbf{38.1^{\pm 0.24}}$ $\mathbf{46.9^{\pm 0.56}}$ $65.3^{\pm 0.44}$ $82.9^{\pm 0.18}$ $\mathbf{45.8^{\pm 0.31}}$ $88.2^{\pm 0.15}$ $\mathbf{65.1^{\pm 0.35}}$ $62.9^{\pm 0.10}$ $\mathbf{63.3^{\pm 0.26}}$
@@
     MoRAM $\mathbf{44.1^{\pm 0.24}}$ $\mathbf{81.6^{\pm 0.34}}$ $\mathbf{64.6^{\pm 0.34}}$ $\mathbf{79.6^{\pm 0.37}}$ $\mathbf{83.9^{\pm 0.36}}$ $84.4^{\pm 0.15}$ $\mathbf{66.5^{\pm 0.24}}$ $89.7^{\pm 0.07}$ $\mathbf{68.4^{\pm 0.38}}$ $64.1^{\pm 0.09}$ $\mathbf{72.7^{\pm 0.17}}$
@@
     MoRAM $\mathbf{37.7^{\pm 0.28}}$ $81.5^{\pm 0.22}$ $\mathbf{70.7^{\pm 0.49}}$ $\mathbf{92.4^{\pm 0.20}}$ $\mathbf{95^{\pm 0.34}}$ $\mathbf{86^{\pm 0.13}}$ $\mathbf{97.6^{\pm 0.19}}$ $92.6^{\pm 0.10}$ $\mathbf{81^{\pm 0.35}}$ $\mathbf{74.7^{\pm 0.06}}$ $\mathbf{80.9^{\pm 0.12}}$
@@
    Method Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
@@
   Zero-shot [63] 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
   LwF [44] 20.0 74.1 79.6 58.1 34.1 48.9 27.7 64.4 15.1 46.9
   WiSE-FT [86] 21.3 79.5 83.3 61.0 39.9 56.5 29.6 68.0 20.8 51.1
   ZSCL [105] 23.0 84.3 87.2 63.0 42.1 65.2 34.6 71.4 40.9 56.9
   MoE-Adapter [96] 17.1 87.2 87.5 58.4 12.6 65.5 35.9 70.0 17.9 50.2
   RAIL-Primal [91] 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
   CoDyRA [50] 23.6 89.2 83 62 51 71.4 38 77.4 39 59.4
   MoRAM 23.6 88.7 83.4 62.6 51.2 69.9 39.3 77.5 39 59.5
@@
   LwF [44] 49.0 27.4 69.7 83.0 65.7 42.2 63.5 33.1 68.5 17.5 52.0
   WiSE-FT [86] 57.9 29.6 77.8 85.4 68.0 51.6 69.3 35.5 71.0 23.0 56.9
   ZSCL [105] 74.4 36.4 86.7 88.7 68.9 50.0 75.1 40.1 72.5 43.7 63.6
   MoE-Adapter [96] 74.4 38.6 87.7 87.3 67.9 50.6 76.5 43.7 72.3 18.8 61.8
   RAIL-Primal [91] 77.9 40.4 85.6 83.3 68.3 62.2 76.6 45.8 80.4 41.7 66.2
   CoDyRA [50] 80 39.2 92.5 85.2 69.2 73.7 79.6 46.2 78.6 44.1 68.8
   MoRAM 80.2 40.1 92.5 84.7 70.1 74 80.1 48.7 78.4 44.4 69.3
@@
   LwF [44] 29.6 17.5 63.0 83.8 67.7 44.9 79.3 44.8 84.6 39.0 55.4
   WiSE-FT [86] 46.1 23.5 71.3 85.7 70.2 59.1 85.5 47.9 82.4 42.8 61.5
   ZSCL [105] 71.7 35.3 86.5 89.2 71.8 52.3 89.8 52.0 77.1 68.4 69.4
   MoE-Adapter [96] 75.1 41.1 87.9 87.1 74.1 89.7 92.6 61.2 81.0 27.4 71.7
   RAIL-Primal [91] 77.7 41.9 86.1 83.3 71.8 91.6 97.3 66.4 94.8 86.9 79.8
   CoDyRA [50] 79 38.6 92.6 86.4 74.7 95.2 93 64.7 81.9 92.2 79.8
   MoRAM 79.3 38.9 93.1 85.4 74.9 96.4 94.1 69.9 82 92.9 80.7
A.9 More comparisons on X-TAIL
To further validate MoRAM ’s robustness, we evaluated it under other continual-learning task orderings, i.e., X-TAIL (Order 2), as shown in Table 17. The results align with those in Table 1, confirming that MoRAM consistently achieves state-of-the-art performance.
@@ A.9 More comparisons on X-TAIL
LWF [44] 129.6 32172
ZSCL [105] 129.6 26290
MoE-Adapters [96] 59.8 22358
CoDyRA [50] 4.4 21770
MoRAM 4.4 21090
A.10 Computation Cost
Table 11 summarizes the per‐task trainable parameters of various continual‐learning methods. Standard LoRA [33], CoDyRA [50], and O-LoRA [80] each introduce $r(d_{\text{in}}+d_{\text{out}})$ new parameters per weight matrix. Mixture‐of‐Experts variants such as MoE-LoRA and MoE-Adapter [96] additionally train a router module to control the usage of each LoRA expert, inducing $d_{\text{in}}$ additional parameters for each expert. LB-CL [61] introduces $r$ additional parameters, mimic the singular values of SVD.
By contrast, MoRAM requires only $rd_{\text{in}}+kd_{\text{out}}$ activated trainable parameters per task, where $k\leq r$ is the activation budget, and the trainable parameter is at most the same as a standard LoRA. Despite the small number of parameters activated and trained, MoRAM achieves superior continual learning performance, and reaches comparable performance in general fine-tuning with only one‐third of the activated parameters of a standard LoRA.
@@ A.10 Computation Cost
LWF [44] and ZSCL [105] perform full‐parameter fine‐tuning, consuming the most parameters and memory. MoE‐Adapters [96] maintains a router with 22 rank‐64 adapter experts (top-2 activated) and a DDAS domain predictor. CoDyRA [50] trains a single rank‐16 LoRA per task, reducing its footprint to 4.4 M parameters. MoRAM introduces 16 rank-1 experts per task, with no additional router, for a total of 4.4 M trainable parameters and keeps a low GPU memory usage, thanks to our novel self‐activated sparse mixture of ranks design.
Table 19: Analysis of post-pruning. After training of each task, we retain the subset required to capture the top 99% of the total activation.
      Method Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
Transfer
     MoRAM 74.5 38.1 46.9 65.3 82.9 45.8 88.2 65.1 62.9 63.3
     MoRAM w/ pruning 75.1 38.0 43.6 68.4 83.9 48.0 88.8 65.3 62.3 63.7
Average
     MoRAM 44.1 81.6 64.6 79.6 83.9 84.4 66.5 89.7 68.4 64.1 72.7
     MoRAM w/ pruning 42.3 81.0 62.9 75.4 82.7 83.2 66.3 89.6 67.6 63.0 71.4
Last
     MoRAM 37.7 81.5 70.7 92.4 95.0 86.0 97.6 92.6 81.0 74.7 80.9
     MoRAM w/ pruning 34.6 80.8 67.9 89.4 94.2 85.3 97.3 92.6 79.8 73.7 79.5
Analysis of post-pruning. While this approach entails linear storage growth, our atomic structure offers a distinct advantage over conventional LoRA. Unlike standard matrices that require complex post-hoc decomposition (e.g., SVD) to compress, our rank-1 experts are independent and can be individually assessed. Consequently, the model inherently supports post-pruning, retaining only atoms that exceed a cumulative activation threshold, allowing for storage reduction with minimal performance degradation if required.
We probe the effects of pruning in Table 19, where we collect the cumulative activation mass of all rank-1 memories during training and retain only the subset required to capture the top 99% of the total activation mass, which on average prunes approximately 30% of the parameter storage. We find that the degradation to performance is minimal, confirming that the model naturally learns a sparse representation where information is concentrated in a concise set of high-utility rank-1 memories.
Figure 4: Averaged activation ratio of each rank-1 expert across three scenarios in Fig. 3
A.11 Extended Visualizations of Aggregated Rank Activations
To complement the qualitative examples in Fig. 3, we include a statistical aggregation of rank-1 expert utilization over the entire test set for each task. This heatmap (Fig. 4) provides a global view of the routing behavior and confirms that the patterns observed in Fig. 3 are representative of the model’s overall dynamics. The heatmap (Fig. 4) visualizes the activation ratio of each rank-1 expert (x-axis) across three scenarios (y-axis) in Fig. 3:
  1. 1.
    Task 1 Data (after Task 1): Consistent with the qualitative results in Fig. 3(a), we observe statistically dominant usage of Rank 0 (airplane semantics) and Rank 11 (background/sky). Rank 11 shows higher overall activation frequency as it captures common background tokens, which constitute a larger portion of image patches than the object itself.
  2. 2.
    Task 1 Data (after Task 2): Crucially, the activation pattern remains virtually unchanged after training on Task 2. The heatmap shows near-zero activation for the newly introduced Task 2 experts (Ranks 16-31). This statistically proves that our routing mechanism is stable: "old" data does not drift to "new" experts, effectively preventing catastrophic forgetting.
  3. 3.
    Task 2 Data (after Task 2): We observe a distinct, dual-mode behavior:
    1. (a)
      Knowledge Reuse: Old Rank 11 is reactivated, confirming that the model reuses the generic "blue sky" feature for the new task.
    2. (b)
      High Diversity: Unlike the concentrated pattern of Task 1 (Aircraft with homogeneous airplane images), Task 2 (Caltech101 with 101 diverse categories) utilizes a broad spectrum of new experts (e.g., Ranks 20, 24, 29, 30). This aligns with our design goal: fine-grained experts allow the model to dedicate different subspaces to the highly diverse semantics of the new task.
We explicitly chose to visualize activations at specific representative layers rather than averaging across the entire depth of the model. In the modern deep models, different layers specialize in distinct feature types (e.g., low-level textures vs. high-level semantics). Averaging expert usage across all layers would smooth out these distinct signatures and obscure the specialized routing behavior we aim to demonstrate.
@@ A.11 Extended Visualizations of Aggregated Rank Activations
(a) Rank Activations of MoRAM on data from Task 1 after learning Task 1.
@@
(b) Rank Activations of MoRAM on data from Task 1 after learning Task 2.
@@
(c) Rank Activations of MoRAM on data from Task 1 after learning Task 10.
@@
(a) Rank Activations of MoRAM on data from Task 1 after learning Task 1.
@@
(b) Rank Activations of MoRAM on data from Task 9 after learning Task 1.
@@
(c) Rank Activations of MoRAM on data from Task 9 after learning Task 9.
@@
A.12 Extended Visualizations of Rank Activations
@@
MoRAM retains task‐specific semantics without forgetting. Fig. 5 shows the activation maps for the same Task 1 image after training on Task 1 (a), Task 2 (b), and the last Task 10 (c). Patches corresponding to the airplane object are outlined in orange. In all three snapshots, Rank at index 0 remains consistently and exclusively activated for those airplane patches, demonstrating that MoRAM has effectively memorized the airplane‐specific knowledge into Rank 0. Even after 10 subsequent tasks, this pattern remains unchanged, indicating that later updates do not overwrite or interfere with the learned airplane representations. In other words, MoRAM effectively memorizes and preserves task‐relevant semantics, thereby mitigating catastrophic forgetting.
MoRAM encodes generic semantics that are reused across tasks. Fig. 6 examines an input image from Task 9 before and after learning Task 9. Panel (a) shows the activation map of data from Task 1 after learning Task 1: Rank 11 (outlined in blue) already responds strongly to sky‐background patches, demonstrating that MoRAM has stored a generic “blue sky” concept in this rank. In panel (b), when we infer on the Task 9 image before training on Task 9, Rank 11 is again activated for the sky regions, confirming that MoRAM reuses this shared knowledge for unseen data. Finally, panel (c) shows the activation map after learning Task 9: Rank 11 remains dedicated to the sky background, while newly initialized ranks specialize in the “car” object semantics. This persistent reuse of Rank 11 across tasks illustrates MoRAM ’s ability to capture and retain common features as reusable memory slots, reducing redundancy and facilitating knowledge reuse.
Figure 7: Statistical analyses on the number of ranks required to capture 99% of the cumulative sum (indicated in red dashed line) of all rank activations. Activations were gathered from the model after training on all tasks, and results are shown for a representative selection of layers and positions within the pre-trained model.
Figure 8: Statistical analyses on the number of ranks required to capture 99% of the cumulative sum (indicated in red dashed line) of all rank activations. Activations were gathered from the model after training on all tasks, and results are shown for a representative selection of layers and positions within the pre-trained model.
@@
A.13 Statistical Analyses of Contributing Rank Activations